OpenStudy (anonymous):

Could someone show me step by step how to integrate: tan (x/2)? I'm not sure if I overlooked any methods.

OpenStudy (zarkon):

\[\tan(x/2)=\frac{\sin(x/2)}{\cos(x/2)}\]

OpenStudy (zarkon):

\[u=\cos(x/2)\]

myininaya (myininaya):

\[\int\limits_{}^{}\frac{\sin(x/2)}{\cos(x/2)}dx, let u=\cos(x/2)=>du=\frac{-1}{2}\sin(x/2) dx\]

myininaya (myininaya):

\[=>-2du=\sin(x/2)dx\]

OpenStudy (zarkon):

you know you can use \text{} in latex to display regular text in math mode. then you can get... \[\int\limits_{}^{}\frac{\sin(x/2)}{\cos(x/2)}dx,\text{ let }u=\cos(x/2)=>du=\frac{-1}{2}\sin(x/2) dx\]

OpenStudy (anonymous):

Thank you very much. I got it now.

myininaya (myininaya):

zarkon so n-spire is the best calculator right?

myininaya (myininaya):

what model of n-spire?

myininaya (myininaya):

if i buy one, what should i make sure comes with it?

myininaya (myininaya):

thanks :)

OpenStudy (zarkon):

awesome calculator

OpenStudy (zarkon):

I don't have the color one (CX)...I just have the black n white n-spire cas

myininaya (myininaya):

someone stole my ti83 so i was thinking about investing a new calculator i love that ti83 too :(

OpenStudy (zarkon):

I have two 83's (83 and 83+)

myininaya (myininaya):

but you know they probably really needed the calculator since they are poor

OpenStudy (zarkon):

two 82's, and 89, 92+ and my nspire

myininaya (myininaya):

so i donated but i don't know who too

myininaya (myininaya):

wow so many

myininaya (myininaya):

you know i'm starting my phd in applied science this semester?

OpenStudy (zarkon):

also have 4 scientific calculators ;)

OpenStudy (zarkon):

no graphinf

OpenStudy (zarkon):

cool

OpenStudy (zarkon):

you have a masters?

myininaya (myininaya):

i only need 2 lecture classes and the rest is research

myininaya (myininaya):

yes

OpenStudy (zarkon):

awesome

myininaya (myininaya):

eventhough i can't do algebra sometimes lol

OpenStudy (zarkon):

you have an adviser in mind?

myininaya (myininaya):

yes

OpenStudy (zarkon):

good...that is really the most important thing to know :)

myininaya (myininaya):

so i'm really scared and excited at the same time

OpenStudy (zarkon):

is there a specific area you are looking to do research in?

OpenStudy (zarkon):

I did mine in bioinformatics

OpenStudy (zarkon):

probabilistic modeling

myininaya (myininaya):

yes i asked everyone in the math department hey can we look into some advanced number theory stuff (i wanted to be an expert in that) and no on was interested because no one had expertise in number theory so i went to the computer department and found someone who does applied cryptography and asked him if he would be interested in theory only and he said yes :) and so i found my partner in crime

OpenStudy (zarkon):

nice

myininaya (myininaya):

theory of cryptography*

OpenStudy (zarkon):

sounds like fun

myininaya (myininaya):

which is everything i want anyways

OpenStudy (anonymous):

wow, color screen

myininaya (myininaya):

number theory group theory set theory algebraic structures abstract algebra

OpenStudy (zarkon):

cool

OpenStudy (zarkon):

65536 colors

myininaya (myininaya):

i didn't get into advanced differential equations

myininaya (myininaya):

also graph theory add that to crypto-list

OpenStudy (zarkon):

I've never really studied cryptography so I don't really know what all is involved.

OpenStudy (zarkon):

sounds like a lot though

myininaya (myininaya):

my goal is to eventually construct my own algorithm for encryption you know like the RSA

OpenStudy (zarkon):

gotcha

myininaya (myininaya):

i need to test it using statitics like see what the probability of it holding its security over a certain time interval

myininaya (myininaya):

since i'm not a programmer

myininaya (myininaya):

math is the only way for me to test it

OpenStudy (zarkon):

ic

myininaya (myininaya):

i don't know how i will do that yet using stat

myininaya (myininaya):

but im sure i will find out how

myininaya (myininaya):

eventually

OpenStudy (zarkon):

I'm sure you will

myininaya (myininaya):

so everyone else is doing finite element and i think thats gross looking lol

OpenStudy (zarkon):

lol

myininaya (myininaya):

you got any number theorist at your college?

OpenStudy (zarkon):

no

myininaya (myininaya):

is that not a popular choice?

OpenStudy (zarkon):

my department is small.

OpenStudy (zarkon):

we only have 2 Ph.D.'s in the department :)

OpenStudy (zarkon):

me- Mathematical Statistics and the other is abstract algebra (group theory)

myininaya (myininaya):

sounds like my school we only have 2 phds as well but one barely teaches because they are an adminstrator (the one i teach at) the one i teach at doesn't have a phd offering of any kind

OpenStudy (zarkon):

it is sort of nice since I get to teach most of the upper level classes

myininaya (myininaya):

i think i would like to teach number theory once i got comfortable with it

OpenStudy (zarkon):

the other prof at my college has taught that class...I have not

OpenStudy (zarkon):

it is interesting stuff

myininaya (myininaya):

yes i agree i took number theory during the summer one semester at UALR and i remember some stuff but i totally need a review my mind doesn't hold very much info anymore :(

myininaya (myininaya):

i want to take memory pills lol

OpenStudy (zarkon):

that would be nice...i need those too

OpenStudy (zarkon):

well...good luck and have fun.

myininaya (myininaya):

later

OpenStudy (zarkon):

later